OpenStudy (anonymous):

the density of an ideal gas a is 1.43g/l at STP. determine the density of a at 290 kelvin and 700 torr(mm)

OpenStudy (cuanchi):

PV = nRT ; n = m/MM ; d= m/V ; P1V1/n1T1 =P2V2/T2n2 Replace n= m/MM in both sides, then replace m/V1 by density 1 and m/v2 by density 2. Isolate d2 d2= d1T1P2/(T2P1)

OpenStudy (cuanchi):

STP: P1=760 tor, T1=273K

OpenStudy (anonymous):

thanx